Green Thumb Gardening is a small gardening service that uses activity-based costing to estimate costs for pricing and other purposes. The proprietor of the company believes that costs are driven primarily by the size of customer lawns, the size of customer garden beds, the distance to travel to customers, and the number of customers. In addition, the costs of maintaining garden beds depends on whether the beds are low maintenance beds (mainly ordinary trees and shrubs) or high maintenance beds (mainly flowers and exotic plants). Accordingly, the company uses the five activity cost pools listed below:

Activity Cost Pool Activity Measure
Caring for lawn Square feet of lawn
Caring for garden beds–low maintenance Square feet of low maintenance beds
Caring for garden beds–high maintenance Square feet of high maintenance beds
Travel to jobs Miles
Customer billing and service Number of customers

The company already has completed its first stage allocations of costs and has summarized its annual costs and activity as follows:

  

Activity Cost Pool Estimated
Overhead
Cost
Expected Activity
Caring for lawn $ 78,600 165,000 square feet of lawn
Caring for garden beds–low maintenance $ 31,200 27,000 square feet of low maintenance beds
Caring for garden beds–high maintenance $ 59,010 21,000 square feet of high maintenance beds
Travel to jobs $ 3,200 18,000 miles
Customer billing and service $ 6,700 22 customers

Required:

Compute the activity rate for each of the activity cost pools.
Answer #1
Activity cost pool Activity rate Formulas
Caring for lawn 0.48 per square ft of lawn =78600/165000
Caring for garden beds-low maintenance 1.16 per square ft of low maintenance beds =31200/27000
Caring for garden beds-high maintenance 2.81 per square ft of high maintenance beds =59010/21000
Travel to jobs 0.18 per mile =3200/18000
Customer billing and service 304.55 per customer =6700/22